# Data-Free Group-Wise Fully Quantized Winograd Convolution via Learnable Scales

### v1

@@ 5.2 Text-to-image generation
W8A8 Winograd F(4,3) 8/8 23.38 30.07
Learned scales F(6,3) 8/8 27.05 29.58
@@
W8A8 Winograd F(4,3) 8/8 28.19 29.57
Learned scales F(6,3) 8/8 34.00 28.93
@@
W8A8 Winograd F(4,3) 8/8 22.29 31.70
Learned scales F(6,3) 8/8 20.72 31.51
@@
W8A8 Winograd F(4,3) 8/8 22.10 31.66
Learned scales F(6,3) 8/8 20.64 31.47
@@
Table 5: ResNets Top-1 accuracy results on the ImageNet dataset.
@@ 7.1 Standard Winograd transforms
$=\begin{bmatrix}[r]1&1&1&1&1&1&1&0\\ 0&1&-1&2&-2&\nicefrac{{1}}{{2}}&-\nicefrac{{1}}{{2}}&0\\ 0&1&1&4&4&\nicefrac{{1}}{{4}}&\nicefrac{{1}}{{4}}&0\\ 0&1&-1&8&-8&\nicefrac{{1}}{{8}}&-\nicefrac{{1}}{{8}}&0\\ 0&1&1&16&16&\nicefrac{{1}}{{16}}&\nicefrac{{1}}{{16}}&0\\ 0&1&1&32&-32&\nicefrac{{1}}{{32}}&-\nicefrac{{1}}{{32}}&0\\ \end{bmatrix}$ (30)
@@
$=\begin{bmatrix}[r]1&0&0\\ -\nicefrac{{2}}{{9}}&-\nicefrac{{2}}{{9}}&\nicefrac{{2}}{{9}}\\ -\nicefrac{{2}}{{9}}&\nicefrac{{2}}{{9}}&\nicefrac{{2}}{{9}}\\ \nicefrac{{1}}{{90}}&\nicefrac{{1}}{{45}}&-\nicefrac{{2}}{{45}}\\ \nicefrac{{1}}{{90}}&-\nicefrac{{1}}{{45}}&\nicefrac{{2}}{{45}}\\ \nicefrac{{32}}{{45}}&\nicefrac{{16}}{{45}}&\nicefrac{{8}}{{45}}\\ \nicefrac{{32}}{{45}}&-\nicefrac{{16}}{{45}}&\nicefrac{{8}}{{45}}\\ 0&0&1\end{bmatrix}$ (32)
@@ 7.2 Learned Winograd scales
W8A8 Winograd F(4,3) 8/8 23.38 30.07
Learned scales F(6,3) 8/8 27.05 29.58
W8A8 Winograd F(4,3) 8/8 228.18 15.17
Learned transforms F(6,3) 8/8 360.14 7.46
@@ 9 Comparison of group-wise quantization against other quantization methods
Table 8: Results on the Stable Diffusion V1.4 model for group-wise quantization, Winograd convolution, AKL autoencoder and DPMSolver++ sampler with 25 steps. Comparison with the results from Q-Diffusion [17].
@@
Table 9: Results on the Stable Diffusion V1.4 model for group-wise quantization, Winograd convolution, AKL autoencoder and PLMS sampler with 25 steps. Comparison with the results from Q-Diffusion [17].
@@ 10 Transferability of learned Winograd scales across datasets
Because Winograd scales are learned from random noise inputs in our method, the same Winograd scale used for one dataset should scale to other datasets also. Data-free finetuning makes Winograd scale values transferable across datasets. Table 10 shows the accuracy results for the CIFAR10 dataset, whereas the same learned Winograd scale values used for the ImageNet dataset previously are applied to another dataset, the CIFAR10 dataset in this case. Winograd, using learned scales, can successfully restore the accuracy of the full-precision ResNet network for both tile sizes. This confirms the transferability of our learned Winograd scales across datasets.
@@
Table 10: ResNets Top-1 accuracy results on the CIFAR-10 dataset.

### v2

@@ 5.2 Text-to-image generation
W8A8 Winograd F(4,3) 8/8 24.51 29.87
Learned scales F(6,3) 8/8 26.58 29.65
@@
W8A8 Winograd F(4,3) 8/8 27.09 29.53
Learned scales F(6,3) 8/8 27.69 29.46
@@
W8A8 Winograd F(4,3) 8/8 19.65 31.63
Learned scales F(6,3) 8/8 20.52 31.53
@@
W8A8 Winograd F(4,3) 8/8 21.47 31.31
Learned scales F(6,3) 8/8 20.56 31.55
Table 5: ResNets Top-1 accuracy results on the ImageNet dataset.
@@ 7.1 Standard Winograd transforms
$=\begin{bmatrix}[r]1&1&1&1&1&1&1&0\\ 0&1&-1&2&-2&\nicefrac{{1}}{{2}}&-\nicefrac{{1}}{{2}}&0\\ 0&1&1&4&4&\nicefrac{{1}}{{4}}&\nicefrac{{1}}{{4}}&0\\ 0&1&-1&8&-8&\nicefrac{{1}}{{8}}&-\nicefrac{{1}}{{8}}&0\\ 0&1&1&16&16&\nicefrac{{1}}{{16}}&\nicefrac{{1}}{{16}}&0\\ 0&1&-1&32&-32&\nicefrac{{1}}{{32}}&-\nicefrac{{1}}{{32}}&0\\ \end{bmatrix}$ (30)
@@
$=\begin{bmatrix}[r]1&0&0\\ -\nicefrac{{2}}{{9}}&-\nicefrac{{2}}{{9}}&-\nicefrac{{2}}{{9}}\\ -\nicefrac{{2}}{{9}}&\nicefrac{{2}}{{9}}&-\nicefrac{{2}}{{9}}\\ \nicefrac{{1}}{{90}}&\nicefrac{{1}}{{45}}&\nicefrac{{2}}{{45}}\\ \nicefrac{{1}}{{90}}&-\nicefrac{{1}}{{45}}&\nicefrac{{2}}{{45}}\\ \nicefrac{{32}}{{45}}&\nicefrac{{16}}{{45}}&\nicefrac{{8}}{{45}}\\ \nicefrac{{32}}{{45}}&-\nicefrac{{16}}{{45}}&\nicefrac{{8}}{{45}}\\ 0&0&1\end{bmatrix}$ (32)
@@ 7.2 Learned Winograd scales
W8A8 Winograd F(4,3) 8/8 24.51 29.87
Learned scales F(6,3) 8/8 26.58 29.65
W8A8 Winograd F(4,3) 8/8 204.00 17.13
Learned transforms F(6,3) 8/8 371.86 3.38
@@ 8 Comparison with learning transformation matrices instead of Winograd scales
9 Advantage of data-free approach over calibration data and other Winograd methods
Using our paradigm, we fine-tuned Winograd scales and transformation matrices in both end-to-end and BRECQ [18] modes. We compute loss using the difference between the generated images or features of the fully group-wise quantized model and its FP16 counterpart. We randomly select 10k prompts from the poloclub/diffusiondb [44] dataset for calibration and use the same setup as when training with random noise. It is more challenging to set up an end-to-end training pipeline, and the diffusion model as a whole needs to be kept on the GPU, hence requiring longer training time and more memory. It can be seen in Table 8 that learning scales with calibration data is also effective for $F(4,3)$, but less so for $F(6,3)$ than learning with noise. The poor accuracy for calibration data when compared to our method may be attributed to the small sample size of 10k prompts, but using more samples may result in overfitting. Similar to data-free training, learning Winograd transforms with calibration data failed to produce good-quality generations. Furthermore, training with BRECQ takes significantly longer (>10×>10$\times$> 10 ×) due to a much more complex training pipeline, with each layer having its own trained Winograd scales, as opposed to a single set of Winograd scales for all layers of one network in our method.
Channel balancing (BQW) [6] only quantizes the Hadamard product, which we could achieve using only group-wise quantization. PAW+FSQ [5] is perhaps not suitable for large-scale diffusion models: transformation matrices are difficult to finetune with the use of calibration data, as shown in Table  8. The use of calibration data for PAW can make it difficult to ensure generalizability to unseen downstream tasks for foundation models. FSQ optimizes hadamard product output (Y) quantization, whereas our learned scales jointly optimize the entire Winograd algorithm, including transformation matrices and Y.
@@ 10 Comparison of group-wise quantization against other quantization methods
Table 8: Train Winograd scales and transforms using calibration data.
IF-0.9B, COCO2017-5k, AKL
Model tile size Bits FID($\downarrow$) CLIP($\uparrow$)
Learned scales F(4,3) 8/8 24.51 29.87
noise F(6,3) 8/8 26.58 29.65
Learned scales F(4,3) 8/8 24.48 29.85
Calibration data E2E F(6,3) 8/8 36.18 29.28
Learned transforms F(4,3) 8/8 138.94 21.05
Calibration data E2E F(6,3) 8/8 262.10 14.20
SD-1.4, COCO2017-5k, AKL, DPMSolver++
FP16 N/A 16/16 21.63 31.72
Learned scales noise F(6,3) 8/8 20.75 31.57
Learned scales BRECQ F(6,3) 8/8 170.24 23.24
Table 9: Results on the Stable Diffusion V1.4 model with group-wise quantization, AKL autoencoder and DPMSolver++ sampler with 25 steps. Comparison with the results from Q-Diffusion [17].
@@
Table 10: Results on the Stable Diffusion V1.4 model with group-wise quantization, AKL autoencoder and PLMS sampler with 25 steps. Comparison with the results from Q-Diffusion [17].
@@ 11 Transferability of learned Winograd scales across datasets
Because Winograd scales are learned from random noise inputs in our method, the same Winograd scale used for one dataset should be transferable to other datasets. Table 11 shows the accuracy results for the CIFAR10 dataset, whereas the same Winograd scale values learned for the ImageNet dataset models previously are applied to the CIFAR10 dataset models directly. Winograd, using learned scales, can successfully restore the accuracy of the full-precision ResNet networks for both tile sizes. This confirms the transferability of our learned Winograd scales across datasets. Table 12 shows the results from Stable Diffusion V1.5 with learned Winograd scales, using DPMSolver++ sampler with 25 sampling steps and a classifier-free guidance scale of 5.0 on the MJHQ dataset. The same learned Winograd scales work for both the COCO and MJHQ datasets.
Table 11: ResNets Top-1 accuracy results on the CIFAR-10 dataset.
@@
Table 12: Results on the Stable Diffusion V1.5 model with group-wise quantization, AKL autoencoder and DPMSolver++ sampler with 25 steps.
SD-1.5, MJHQ-2k, AKL, DPMSolver++
Model tile size Bits FID($\downarrow$) CLIP($\uparrow$)
FP16 N/A 16/16 41.40 25.78
Learned scales F(4,3) 8/8 44.03 25.53
noise F(6,3) 8/8 45.56 25.56
@@ 12 Comparison to prior QAT works for fully quantizing Winograd
13 Highly optimized kernels design for text-to-image generation inference on CPUs
While C/C++ runtimes like stablediffusion.cpp [1] demonstrate performance and potential on CPUs, the baseline group quantized kernels have significant compute overheads (the leftmost bar in Figure 6(a)). As a result, a higher proportion of compute instructions do not perform multiplies, i.e., real work, rendering them unsuitable for meeting the required latency requirements for text-to-image generation model variants deployed on commodity CPUs.
To reduce compute overhead and increase the reuse of the input matrices, as well as the use of MAC and vector operations, our highly optimized matrix multiply kernels consider a series of consecutive rows and columns from the two input matrices at once. The use of multiple rows and columns allows for greater reuse of quantized matrices and associated scale factors, as well as fewer load operations. Furthermore, to reduce the high overhead from reduction operations and quantized operand unpacking operations in group-quantized matrix multiply kernels, as well as improve MAC unit utilization and thus the percentage of useful work, our highly optimized kernels perform the following optimizations:
  • Amortize the cost of loading operands and the cost of weight unpacking across multiple output channels, resulting in fewer load operations, increased operand reuse, and greater use of vector operations.
  • Eliminate the overhead of frequently occurring explicit reduction operations in the matrix multiply kernel (required to accumulate partial dot products from different vector lanes of SIMD fused multiply-accumulate (dot product) operations to obtain the final dot product result for a quantized weight group) of group-wise quantized kernels by ensuring that different vector lanes of dot product instructions operate on weight elements and groups from different output channels so that dot product instructions can take advantage of their implicit accumulate (reduction) operations.
  • Solve the resulting non-contiguous access patterns of quantized groups across channels in memory by reformatting weights offline to match the compute order and packing, then storing them in memory in reordered format.
  • Maximize the use of the available high MAC throughput matrix-matrix multiply-accumulate instruction, which can perform twice as many MAC operations when compared to an equivalent SIMD dot product instruction.
The benefits of our matrix multiply kernel optimizations for group-wise quantized Winograd convolution and attention layers for Arm CPUs will easily extend to GPUs, neural engines, and x86 processors for the following reasons: (1) Our kernels reduce the number of instructions in the critical path of the group-wise quantized kernels through efficient aligning and reordering of operands of group-quantized formats to take advantage of implicit reduction operations of vector dot product instructions, increase operand reuse, maximize MAC unit utilization, minimize overhead and memory accesses, (2) The vector and matrix-matrix multiply operations used in optimized group-wise quantized kernels are available in other existing CPU and GPU architectures, so there is no need to add new instructions to the CPU/GPU ISA.